# Hope vs. Hate: Understanding User Interactions with LGBTQ+ News Content in Mainstream US News Media through the Lens of Hope Speech

### v1

@@ Abstract
This paper makes three contributions. First, via a substantial corpus of 1,419,047 comments posted on 3,161 YouTube news videos of major US cable news outlets, we analyze how users engage with LGBTQ+ news content. Our analyses focus both on positive and negative content. In particular, we construct a fine-grained hope speech classifier that detects positive (hope speech), negative, neutral, and irrelevant content. Second, in consultation with a public health expert specializing on LGBTQ+ health, we conduct an annotation study with a balanced and diverse political representation and release a dataset of 3,750 instances with fine-grained labels and detailed annotator demographic information. Finally, beyond providing a vital resource for the LGBTQ+ community, our annotation study and subsequent in-the-wild assessments reveal (1) strong association between rater political beliefs and how they rate content relevant to a marginalized community; (2) models trained on individual political beliefs exhibit considerable in-the-wild disagreement; and (3) zero-shot large language models (LLMs) align more with liberal raters.
@@ 1 Introduction
From crowdsourced storymapping projects sharing stories of love, loss, and a sense of belonging [1] to safe, anonymous spaces to seek resources [2] and dating platforms [3] – the internet and modern technologies play a positive role in health and well-being of the LGBTQ+ community in various ways. However, cyberbullying [4], exposure to dehumanization through news media [5], and more recently, worrisome homophobic biases in large language models [6] - are some of the modern technology perils the community grapples with.
How do mainstream US cable news outlets cover the LGBTQ+ community and how do users interact with such news content? Via a substantial corpus of more than 130 million YouTube comments on more than 300K YouTube news videos uploaded on three major US cable news outlets, this paper investigates how LGBTQ+ discussions are situated in mainstream US political discourse. While prior literature has analyzed LGBTQ+ discourse in print media through the lens of a dehumanization framework [5], such analyses consulted a single source of news (The New York Times) and did not consider reader responses to LGBTQ+ news. In this work, we consider three major US cable news networks representing diverse political views (Fox News, CNN, and MSNBC) with a key goal to understand how users interact and engage with LGBTQ+ news content. In doing so, we not only focus on identifying negative discussions about the LGBTQ+ community, one of our key contributions is operationalizing the detection of positive, hope speech championing the disadvantaged minority in the broader spirit of counter-speech literature [7, 8, 9, 10, 11].
@@
This is why we have 500+ million guns in america. We need to start using them. If a man walked in my daughters bathroom, that would be his last day on earth. People arent gonna keep tolerating this. Theres about to be rainbow street justice youre gonna start seeing. When comes to my, or anyones kids, and i frankly dont care what your laws. Never did in the first place, but when it comes to my kid……ill wake up and chose violence EVERY time. You christians can worry about the jesus nonsense. Keep “turning the other cheek” and here tf we are. The right republicans and media love their witty banter BUT WHAT EXACTLY TF DO YOU ACTUALLY DO?!!!!!!!!!!!!! NOTHING. Negative
[…]
LISTEN PEOPLE: EVEN IF YOU DONT SUPPORT IT, YOU CAN RESPECT US AT LEAST. ALL WE WANT IS BASIC HUMAN DECENCY. LITERALLY JUST THAT. me and my girlfriend got spat at on a bus once, and then called homophobic slurs. we just want respect, being gay doesn’t change the fact that I’m a human. JUST BE NICE TO EVERYONE AND STOP CARING SO MUCH ABT OTHER PEOPLE AND THEIR BUSINESS. Positive
Reading these comments are making me cry. I’m 13 and my Dad is transgender. 99% of Doctors accept it. Why are you one to argue with Science? Please, if you are a bigot, don’t reply to this comment. It will make me cry. I love my Dad, and you ignorant people who don’t accept science won’t change that. Please don’t reply to this comment unless you have something nice to say. Positive
Table 1: Illustrative examples of positive and negative YouTube comments identified by our classifier in the wild
While hate speech detection and mitigation strategies have been extensively researched (see, e.g., [12]), relatively little attention has gone into the opposite: hope speech. Defined as comments and posts that inspire optimism and diffuse hostility in online spaces [9, 10], hope speech is crucial for groups who face a disproportionate amount of hate speech. In the current US cultural climate, where issues of race, gender, and sexuality are at the forefront of the so-called “culture war” [13], there is a need for tools that can detect both positive and negative content as listed in Table 1 and increase understanding of relevant online conversations.
@@
(1). In consultation with a health expert specializing in LGBTQ+ health for more than two decades, we curate a novel dataset of 3,750 instances with fine-grained labels: neutral, irrelevant, positive (hope speech), and negative [^1] Will be made publicly available upon acceptance.. Each instance is labeled by three annotators (one Republican, one Democrat, and one independent) ensuring diverse and balanced political perspectives in the annotation process. Several raters self-identified as being part of the LGBTQ+ community. (2). We analyze the association between rater political beliefs and how they rate content relevant to LGBTQ+ community. (3) We analyze the alignment of large language models with political beliefs in connection with LGBTQ+ content and investigate how rater biases perpetuate into fine-tuned models. (4). We provide novel insights into audience engagement patterns and sentiment dynamics surrounding LGBTQ+ content, enhancing our understanding of LGBTQ+ discussions within mainstream US political discourse. To our knowledge, no such study exists at our scale.
@@
2 Related Work
Our work is closely related with the hope speech literature [9, 14, 10] and the broader literature of counter speech [7, 8, 11, 15, 16]. Our work contributes to this literature through our unique focus on LGBTQ+ discourse in mainstream US politics and our investigation on how rater subjectivity and political leanings.
Political polarization in the US is widely studied across multiple disciplines in rich and diverse settings that include congressional voting patterns on policy issues [17], mate selection [18], allocation of scholarship funds [19], and annotating offensive content [20, 21]. While how political leanings might associate with the task of rating negative content has been studied both in the context of US politics [20, 21] and on hot button issues such as reproductive rights, gun control/rights [21], and US infrastructure [22], our work demonstrates that political identities of the raters may influence how they perceive positive content for the LGBTQ+ community. Our study thus contributes to the broader literature of annotator subjectivity [23, 24, 25, 26, 20, 27, 28, 21, 29, 30].
Our annotation study design is grounded in prior literature [20, 21] and draws from [21] in seeking diverse and balanced political perspectives including from the independents. Our work also touches upon political bias in LLMs as we observe that zero-shot classification of LGBTQ+ content of several models are more aligned with liberal raters [31, 32]. To tackle class imbalance, we construct an ensemble active learning pipeline much akin to [14] leveraging well-known active learning strategies (e.g., certainty sampling [33] and margin sampling [34]).
@@ 3 Dataset and Annotation Study Design
Figure 1 presents a schematic diagram of our data collection process described in details next.
@@ 3.1 YouTube News Videos
We consider a dataset of 333,032 YouTube videos uploaded between 04/10/2008 to 08/31/2024 by the official YouTube channels of three major US cable news networks: Fox News, CNN, and MSNBC. First introduced in [35], this dataset represents a reliable snapshot of US political discourse encompassing diverse political perspectives over a substantial period of time. This dataset has found prior use in studying (1) election misinformation [36]; (2) health misinformation [37]; and (3) annotation subjectivity [21].
@@ 3.2 Identifying LGBTQ+ Relevant Videos
In contrast to prior literature predating highly capable LLMs, which used keyword filtering to prune large datasets for relevance (see, e.g., [38, 39, 40]), we conduct our filtering using LLMs using a two-step process. First, for each video, we pass the video title and description to GPT-4o mini prompting it to classify the video as specifcially relevant to the LGBTQ+ community or not (Appendix contains prompt details in § B). This step prunes our initial set of 333,032 videos down to 5,435 videos. Next, for each of these 5,435 videos, we use the same prompt and pass the video title and description to GPT-4o. We use this two-step process because although the GPT-4o model performed better and ended up with slightly fewer false positives, it was considerably more expensive to run it on all 333,032 videos. Overall, through this two-step process, we obtain 3,161 relevant videos (denoted by $\mathcal{V}_{\textit{LGBTQ+}}$). A manual inspection of 300 random videos (100 from each news outlet) reveals only 2 false positives indicating that $\mathcal{V}_{\textit{LGBTQ+}}$ is highly relevant for LGBTQ+ issues.
Video title Video description GPT-4o explanation Classification
Table 2: Videos identified as LGBTQ+ relevant by our GPT-4 pipeline
Table 2 shows an example true positive and a false positive found in our results. The first shows a case in which the video ended up not being LGBTQ+ relevant, except for the fact that the person featured is a director of trans care, but the video does not focus on that fact. It is interesting that GPT-4o actually acknowledges this in its explanation, however did end up still flagging it as relevant. The second is an example that identifies the bathroom bill as relevant to LGBTQ+ issues.
@@ 3.3 Constructing $\mathcal{D}_{\textit{hope}}$
For each video in $\mathcal{V}_{\textit{LGBTQ+}}$, we collect user comments using publicly available YouTube API yielding 1,419,047 comments. From these comments, we construct $\mathcal{D}_{\textit{hope}}$, a dataset with fine-grained labels of posts supportive of the LGBTQ+ community. In consultation with the health expert specializing on LGBTQ+ health research for more than two decades, we identify four broad categories: positive (hope speech); negative; irrelevant; and neutral. Detailed annotation guidelines are present in the Appendix (§ C).
@@
    We first construct a seed set of 1,950 instances using a collaborative human-LLM annotation framework. This seed set is annotated by crowd-sourced annotators for the aforementioned four categories (details are described in § F.0.1). This step yields $\mathcal{D}_{\textit{seed}}$ consisting of 649 positives, 150 negatives, 87 neutrals, and 660 irrelevants and a validation set $\mathcal{D}_{\textit{eval}}$ of 250 instances set aside for all performance evaluation.
@@
    Next, we conduct minority class certain sampling [33] of additional 900 instances to address class imbalance (details are described in § F.0.2). This step yields $\mathcal{D}_{\textit{certainty}}$ consisting of 680 positives, 682 negatives, 227 neutrals, and 764 irrelevants.
@@
    Finally, we conduct margin sampling [34], a well-known variant of uncertainty sampling (described in § F.0.2). This step yields $\mathcal{D}_{\textit{margin}}$ consisting of 824 positives, 947 negatives, 314 neutrals, and 1,020 irrelevants.
@@ 3.3.1 Annotating $\mathcal{D}_{\textit{hope}}$
We use Prolific, a well-known crowd-sourcing platform, to gather and pay raters and hosted the survey on Qualtrics. Our study follows a similar annotation design as [21] in the sense that we consider raters evenly distributed across three political leanings (Republican, Democrat, and Independent). We split our corpus into a set of 65 batches, each containing an equal amount of comments from each channel and containing/not containing hope speech. Each batch is labeled by three annotators (one Republican, one Democrat, and one Independent). Since we do not want to expose raters to objectionable content for a sustained period, following standard practices, our batch size is limited to 30 instances. We compensate each rater \$0.10 for each example. This yields a compensation of \$3 for each batch. We observe a median task completion time of 11:53 minutes, implying an hourly wage estimate of \$15.16, more than the minimum wage. This annotation compensation is grounded in prior literature [41, 42, 21]. We also left an unstructured text feedback input for each of the workers. No one complained about the compensation. Many subjects thanked us for providing an interesting and thought-provoking task, while some responded to the survey with homophobia, which might lead to interesting follow-on research (Appendix contains sample responses). Our study was reviewed by our Institutional Review Board and was deemed as exempt.
Our study asked the raters to annotate each comment with one of the following labels (in regards to sentiment towards the LGBTQ+ community) (1) Positive, (2) Negative, (3) Neutral, or (4) Irrelevant. In pilot test surveys, we tried asking raters to pick one of only two labels, Positive or Not Positive. This appeared to make the task more challenging as due to less nuance and granularity in the annotation process, shown by lower agreement between annotators. The choice to switch to four labels (advised by the health expert) also benefited us by introducing more detailed information for analysis and training of models.
On our final annotated dataset, the Fleiss’ $\kappa$ for all four labels is 0.431 and for just two labels (HopeSpeech (positive) and not-HopeSpeech (neutral, negative, and irrelevant)) is 0.485. This shows relatively moderate agreement between all annotators. For a challenging task like this, involving analysis of sentiment towards a specific community, which can be affected by personal beliefs and values, we observe considerably high agreement. Interestingly, the agreement remains relatively stable when expanding from two to four labels, indicating consistent annotator judgment across different levels of classification detail.

### v2

@@ Abstract
This paper makes three contributions. First, via a substantial corpus of 1,419,047 comments posted on 3,161 YouTube news videos of major US cable news outlets, we analyze how users engage with LGBTQ+ news content. Our analyses focus both on positive and negative content. In particular, we construct a hope speech classifier that detects positive (hope speech), negative, neutral, and irrelevant content. Second, in consultation with a public health expert specializing on LGBTQ+ health, we conduct an annotation study with a balanced and diverse political representation and release a dataset of 3,750 instances with crowd-sourced labels and detailed annotator demographic information. Finally, beyond providing a vital resource for the LGBTQ+ community, our annotation study and subsequent in-the-wild assessments reveal (1) strong association between rater political beliefs and how they rate content relevant to a marginalized community, (2) models trained on individual political beliefs exhibit considerable in-the-wild disagreement, and (3) zero-shot large language models (LLMs) align more with liberal raters.
Trigger Warning: this paper contains offensive material that some may find upsetting.
Hope vs. Hate: Understanding User Interactions with LGBTQ+ News Content in Mainstream US News Media through the Lens of Hope Speech
Jonathan Pofcher   Christopher M. Homan Randall L. Sell     Ashiqur R. KhudaBukhshthanks:   Ashiqur R. KhudaBukhsh is the corresponding author. Rochester Institute of Technology Drexel University jep7453@rit.edu, cmhvcs@rit.edu, randy@drexel.edu, axkvse@rit.edu
@@ 1 Introduction
From crowdsourced storymapping projects sharing stories of love, loss, and a sense of belonging Kirby et al. (2021) to safe, anonymous spaces to seek resources McInroy et al. (2019) and dating platforms Blackwell et al. (2015) – the internet and modern technologies play a positive role in the health and well-being of the LGBTQ+ community in various ways. However, cyberbullying Abreu and Kenny (2018), exposure to dehumanization through news media Mendelsohn et al. (2020), and more recently, homophobic biases in large language models (LLMs) Dutta et al. (2024a, 2025a) - are some of the modern technology perils the community grapples with.
How do mainstream US cable news outlets cover the LGBTQ+ community and how do users interact with such news content? Via a substantial corpus of more than 130 million YouTube comments on more than 300K YouTube news videos uploaded on three major US cable news outlets, this paper investigates how LGBTQ+ discussions are situated in mainstream US political discourse. While prior literature has analyzed LGBTQ+ discourse in print media through the lens of a dehumanization framework Mendelsohn et al. (2020), such analyses consulted a single source of news (The New York Times) and did not consider reader responses to LGBTQ+ news. In this work, we consider three major US cable news networks representing diverse political views (Fox News, CNN, and MSNBC) with a key goal to understand how users interact and engage with LGBTQ+ news content. In doing so, we not only focus on identifying negative discussions about the LGBTQ+ community, one of our key contributions is operationalizing the detection of hope speech championing the disadvantaged minority in the broader spirit of counter-speech literature Benesch (2014); Mathew et al. (2019); Palakodety et al. (2020b); Chakravarthi (2020); Hengle et al. (2024).
@@
This is why we have 500+ million guns in america. We need to start using them. If a man walked in my daughters bathroom, that would be his last day on earth. People arent gonna keep tolerating this. Theres about to be rainbow street justice youre gonna start seeing. When comes to my, or anyones kids, and i frankly dont care what your laws. Never did in the first place, but when it comes to my kid……ill wake up and chose violence EVERY time… Negative
Hey if using the term fg is hate speech, then I stand guilty as charged. I hate what they stand for and I hate the corruption they’re bringing into society, and I especially hate that they are preying on the minds of children in an attempt to make them believe that their sick, perverted lifestyle is normal… Negative
LISTEN PEOPLE: EVEN IF YOU DONT SUPPORT IT, YOU CAN RESPECT US AT LEAST. ALL WE WANT IS BASIC HUMAN DECENCY. LITERALLY JUST THAT. me and my girlfriend got spat at on a bus once, and then called homophobic slurs. we just want respect, being gay doesn’t change the fact that I’m a human… Positive
Reading these comments are making me cry. I’m 13 and my Dad is transgender. 99% of Doctors accept it. Why are you one to argue with Science? Please, if you are a bigot, don’t reply to this comment. It will make me cry. I love my Dad, and you ignorant people who don’t accept science won’t change that…. Positive
Table 1: Illustrative examples of positive and negative YouTube comments identified by our classifier in the wild.
While hate speech detection and mitigation strategies have been extensively researched (see, e.g., Fortuna and Nunes, 2018), relatively little attention has gone into the opposite: hope speech. Defined as comments and posts that inspire optimism and diffuse hostility in online spaces Palakodety et al. (2020b); Chakravarthi (2020), hope speech is crucial for groups who face a disproportionate amount of hate speech. In the current US cultural climate, where issues of race, gender, and sexuality are central to public debate Hartman (2019), there is a need for tools that can detect both hope speech and negative content as listed in Table 1 and increase understanding of relevant online conversations. According to the FBI, one in five hate crimes in the US targets the LGBTQ+ community Federal Bureau of Investigation (2024), and globally, homosexual activity remains punishable by death in six UN member states while only 37 countries recognize same-sex marriage Mendos and Rohaizad (2024). Understanding and improving online discourse around LGBTQ+ issues is thus a vital step toward protecting a vulnerable community.
@@
(1) In consultation with a health expert specializing in LGBTQ+ health for more than two decades, we curate a novel dataset of 3,750 instances with labels: neutral, irrelevant, positive (hope speech), and negative [^1] The dataset is available at https://github.com/Social-Insights-Lab/LGBTQplus_HopeSpeech/.. Each instance is labeled by three raters (one Republican, one Democrat, and one independent) ensuring diverse and balanced political perspectives. 98 raters self-identified as being part of the LGBTQ+ community.
(2) We analyze the association between rater political beliefs and how they rate content relevant to the LGBTQ+ community.
(3) We analyze the alignment of LLMs with political beliefs in connection with LGBTQ+ content and how rater biases affect model fine-tuning.
(4) We provide novel insights into audience engagement patterns surrounding LGBTQ+ content, enhancing our understanding of LGBTQ+ discussions within mainstream US political discourse. To our knowledge, no such study exists at our scale.
2 Related Work
Our work is closely related with the hope speech literature Palakodety et al. (2020b, a); Chakravarthi (2020); Yoo et al. (2021) and the broader literature of counter speech Benesch (2014); Mathew et al. (2019); Hengle et al. (2024); Saha et al. (2022); Gupta et al. (2023); Saha et al. (2024). Our work contributes to this literature through our unique focus on LGBTQ+ discourse in mainstream US politics and our investigation on rater subjectivity and political leanings.
Political polarization in the US is widely studied across multiple disciplines in rich and diverse settings that include congressional voting patterns on policy issues Poole and Rosenthal (1984), mate selection Huber and Malhotra (2017), allocation of scholarship funds Iyengar and Westwood (2015), and annotating offensive content Sap et al. (2022); Weerasooriya et al. (2023). Prior research showed systematic differences in offensive speech annotation based on annotators’ beliefs Sap et al. (2022); Weerasooriya et al. (2023) and experiences Patton et al. (2019). While associations between political leanings and annotation of negative content has been studied in US politics Sap et al. (2022); Weerasooriya et al. (2023) and hot-button issues like reproductive rights and gun control Weerasooriya et al. (2023), our work extends this in two ways: first by examining perceptions of positive rather than negative content, and second by focusing specifically on LGBTQ+ discourse. Our study thus contributes to the broader literature of annotator subjectivity Pavlick and Kwiatkowski (2019); Sap et al. (2019); Al Kuwatly et al. (2020); Larimore et al. (2021); Sap et al. (2022); Goyal et al. (2022); Pei and Jurgens (2023); Weerasooriya et al. (2023); Homan et al. (2024); Prabhakaran et al. (2024).
Our annotation study design is grounded in the prior literature Sap et al. (2022); Weerasooriya et al. (2023) and draws from Weerasooriya et al., 2023 and Crowl et al., 2025 in seeking diverse and balanced political perspectives including from the independents. Our work also touches upon political bias in LLMs as we observe that zero-shot classification of LGBTQ+ content of several models are more aligned with liberal raters Feng et al. (2023); Bang et al. (2024). To address class imbalance, we construct an ensemble active learning pipeline much akin to Palakodety et al., 2020a and Khorramrouz et al., 2023 leveraging well-known active learning strategies (e.g., certainty sampling Sindhwani et al. (2009) and margin sampling Scheffer et al. (2001)).
@@ 3 Dataset and Annotation Study Design
Figure 1 presents a schematic diagram of our data collection process. which will be described in more detail in the following section.
@@ 3.1 YouTube News Videos
We consider a dataset of 333,032 YouTube videos uploaded between 04/10/2008 to 08/31/2024 by the official YouTube channels of three major US cable news networks: Fox News, CNN, and MSNBC. First introduced in KhudaBukhsh et al., 2021, this dataset represents a reliable snapshot of US political discourse encompassing diverse political perspectives over a substantial period of time. This dataset has found prior use in studying election-related discourse KhudaBukhsh et al. (2022); Mittal et al. (2024), health-related discourse  Yoo and KhudaBukhsh (2023), and rater subjectivity Weerasooriya et al. (2023); Pandita et al. (2024); Dutta et al. (2025b). We extend this dataset using the same data collection method to include videos up to August 2024.
@@ 3.2 Identifying LGBTQ+ Relevant Videos
We prune our initial set of videos using a two-step process. First, for each video, we pass the video title and description to GPT-4o mini prompting it to classify the video as specifcially relevant to the LGBTQ+ community or not (Appendix contains prompt details in § D). This step prunes our initial set of 333,032 videos down to 5,435 videos. Next, for each of these 5,435 videos, we use the same prompt and pass the video title and description to GPT-4o. While the GPT-4o model performed better and ended up with fewer false positives, we use this two-step process because it was considerably more expensive to run GPT-4o on all 333,032 videos. Overall, we obtain 3,161 relevant videos, denoted $\mathcal{V}_{\textit{LGBTQ+}}$. A random sample of 300 videos from this pipeline (100 from each channel) yielded 2 false positives for related and no false negatives for unrelated, indicating that our pruning steps are robust [^2] Appendix contains examples (Table A18) of LLM explanations for a true positive and a false positive..
Dataset Size Sampling Strategy Positive Negative Neutral Irrelevant
$\mathcal{D}_{\textit{seed}}$ 1,950 Initial crowd-sourced dataset 649 (33.3%) 150 (7.7%) 87 (4.5%) 660 (33.8%)
$\mathcal{D}_{\textit{certainty}}$ 2,850 Certainty sampling 680 (23.9%) 682 (23.9%) 227 (8.0%) 764 (26.8%)
$\mathcal{D}_{\textit{margin}}$ 3,750 Margin sampling Scheffer et al. (2001), a well-known variant of uncertainty sampling 824 (22.0%) 947 (25.3%) 314 (8.4%) 1,020 (27.2%)
$\mathcal{D}_{\textit{eval}}$ 250 Random set of comments taken from $\mathcal{D}_{\textit{seed}}$ held out from training for evaluation 824 (22.0%) 947 (25.3%) 314 (8.4%) 1,020 (27.2%)
Table 2: Evolution of $\mathcal{D}_{\textit{hope}}$ through active learning stages, showing dataset size, sampling methodology, and label distribution.
@@ 3.3 Constructing $\mathcal{D}_{\textit{hope}}$
For each video in $\mathcal{V}_{\textit{LGBTQ+}}$, we collect user comments using publicly available YouTube API yielding 1,419,047 comments. From these comments, we construct $\mathcal{D}_{\textit{hope}}$, a dataset that categorizes content based on its stance toward the LGBTQ+ community. In consultation with the health expert specializing on LGBTQ+ health research for more than two decades, we identify four broad categories: positive (hope speech); negative; irrelevant; and neutral. We define hope speech as content that expresses support, advocacy, or acceptance for the LGBTQ+ community. Negative content is defined as content that expresses opposition, discrimination, or hostility towards the LGBTQ+ community. Neutral content is defined as content that is relevant to LGBTQ+ issues but does not belong to any of the previous two categories (hope speech and negative). Irrelevant content is defined as content that is not related to LGBTQ+ issues. Detailed annotation guidelines are present in the Appendix (§ E).
@@
    We first construct a seed set of 1,950 instances using a collaborative human-LLM annotation framework. This seed set is annotated by crowd-sourced annotators for the aforementioned four categories (details are described in § A.1). This step yields $\mathcal{D}_{\textit{seed}}$ consisting of 649 positives, 150 negatives, 87 neutrals, and 660 irrelevants and a validation set $\mathcal{D}_{\textit{eval}}$ of 250 instances set aside for all performance evaluation.
@@
    Next, we conduct minority class certain sampling Sindhwani et al. (2009) of additional 900 instances to address class imbalance (details are described in § A.2). This step yields $\mathcal{D}_{\textit{certainty}}$ consisting of 680 positives, 682 negatives, 227 neutrals, and 764 irrelevants.
@@
    Finally, we conduct margin sampling Scheffer et al. (2001) (described in § A.2) that yields $\mathcal{D}_{\textit{margin}}$ consisting of 824 positives, 947 negatives, 314 neutrals, and 1,020 irrelevants.
@@ 3.3.1 Annotating $\mathcal{D}_{\textit{hope}}$
We use Prolific for our annotation study, and hosted the survey on Qualtrics. Our study was reviewed by our Institutional Review Board and was deemed as exempt. The annotation study follows a similar annotation design as Weerasooriya et al., 2023 and Crowl et al., 2025 in considering raters evenly distributed across three political leanings (Republican, Democrat, and Independent). We split our corpus into a set of 65 batches, each containing an equal amount of comments from each channel. Each batch is labeled by three annotators (one Republican, one Democrat, and one Independent). Since we do not want to expose raters to objectionable content for a sustained period, following standard practices, we limit our batch size to 30 instances. We also left an unstructured text feedback input for each of the workers. While many raters thanked us for providing an interesting and thought-provoking task, some responded to the survey with homophobia (Appendix contains examples in § I) which might lead to interesting follow-on research.
On our final annotated dataset, we achieve moderate inter-rater agreement. The Fleiss’ $\kappa$ for all four labels is 0.431 and for just two labels (HopeSpeech (positive) and not-HopeSpeech (neutral, negative, and irrelevant)) is 0.485. For a potentially subjective task like ours this agreement is in line with extensive prior literature. For instance, Guest et al., 2021 reported Fleiss’ $\kappa$ of 0.48 while Sanguinetti et al., 2018 reported category-wise $\kappa$ = 0.37. We further note that our observed inter-rater agreement is higher than Gomez et al., 2019 ($\kappa$ = 0.15), Fortuna et al., 2019 ($\kappa$ = 0.17) and Weerasooriya et al., 2023 ($\kappa$ = 0.39). Weerasooriya et al., 2023 studied political subjectivity in offensive content annotation and is closest to our study.
Llama Dubey et al. (2024) Mistral Jiang et al. (2023)
Accuracy 0.5520 0.8480 0.8400 0.8480 0.6640 0.8120 0.8560 0.8480
Macro-averaged metrics:
Precision 0.5468 0.8025 0.7204 0.7993 0.5654 0.6297 0.7947 0.7960
Recall 0.5547 0.6366 0.6866 0.6947 0.5488 0.6001 0.6600 0.6843
F1 Score 0.4768 0.6680 0.6995 0.7162 0.5375 0.6017 0.6880 0.7087
Class-specific F1 scores:
Positive 0.6845 0.8703 0.8727 0.8661 0.8312 0.8448 0.8851 0.8722
Negative 0.3009 0.5806 0.5789 0.6842 0.4545 0.5714 0.5882 0.6486
Neutral 0.2553 0.2759 0.4103 0.3871 0.1765 0.0690 0.3333 0.3871
Irrelevant 0.6667 0.9453 0.9360 0.9275 0.6879 0.9216 0.9453 0.9268
Table 3: Comparison of Llama and Mistral model results on aggregate labels over five independent training runs on $\mathcal{D}_{\textit{eval}}$. $\mathcal{M}(.)$ denotes the zero-shot performance. $\mathcal{M}(\mathcal{D})$ denotes model performance when fine-tuned on $\mathcal{D}$. Table 2 summarizes dataset statistics.